Mike asks Billy to help him run his lemonade stand. He’ll either pay him 2/5 of ½ of the stand’s profits, or 1/3 of ¾ of the stand’s profits. Which should Billy choose to maximize the amount he will earn, and by what amount will his earnings be increased?
A. Billy should choose 1/3 of ¾; he’ll earn 1/15 more.
B. Billy should choose 1/3 of ¾; he’ll earn 1/20 more.
C. Billy should choose 2/5 of ½; he’ll earn 1/15 more.
D. Billy should choose 2/5 of ½; he’ll earn 1/20 more.
E. It doesn’t matter; either way, Billy will earn exactly the same amount.
Please sure me how you get the answer

Answer by checkley77(12844)

2/5*1/2=2/10=.20
1/3*3/4=3/12=.25 THE BETTER DEAL
3/12-2/10=30/120-24/120=6/120=1/20 ANS.B